# Apparent Selection Pressure for Channel Capacity in Bacterial Chemotactic Sensors

### v2

@@ Abstract
Bacterial chemotactic sensing converts noisy chemical signals into running and tumbling. We analyze the static sensing limits of mixed Tar/Tsr chemoreceptor clusters in individual Escherichia coli cells using a heterogeneous Monod–Wyman–Changeux (MWC) model. By sweeping a seven-dimensional parameter space, we compute three sensing performance metrics—channel capacity, effective Hill coefficient, and dynamic range. Across E. coli–like parameter regimes, we consistently observe pronounced local maxima of channel capacity, whereas neither the effective Hill coefficient nor the dynamic range exhibit comparable optimization. The capacity-achieving input distribution is bimodal, which implies that individual cells maximize information by sampling both low- and high concentration regimes. Together, these results suggest that, at the individual-cell level, channel capacity may be selected for in E. coli receptor clusters.
@@ I Introduction
A model organism of study for this are bacteria. Their bacterial chemotactic receptors, responsible for the movement of bacteria cells in response to environmental chemical stimuli, have been the subject of a number of beautiful models [11, 19, 13, 14]. In order to survive, E. coli cells must take in information about chemoattractant concentration [2] so as to move to maximize chemoattractant intake. Recent measurements in E. coli cells have demonstrated that they operate close to optimal information transfer for gradient-climbing [12]. A recent study showed that bacteria optimize predictive information transfer given resource constraints [17]. Furthermore, basic theoretical analyses say that populations of bacteria should maximize predictive information subject to constraints on resources [10].
However, no studies have addressed the channel capacities of individual bacterium. While some studies on E. coli chemotactic receptors focus on the population level and dynamical behavior of cells [9, 10], the present study provides an analysis of the chemotactic ability of a single E. coli cell. Specifically, the mixed Tar/Tsr chemoreceptor clusters within the E. coli cell are modeled with a heterogeneous Monod-Wyman-Changeux molecule [11] under the assumption of adaptation to negligible background chemoattractant concentration [4]. Three sensor performance metrics (channel capacity, effective Hill coefficient, and dynamic range) are obtained by sweeping a seven-dimensional parameter space. By locating literature-based strains in these seven-dimensional landscapes, we find that channel capacity appears to have been driven to a local maximum for all available strains. Dynamic range and effective Hill coefficient are positively correlated with channel capacity across the seven-dimensional parameter space, but upon closer examination, do not appear to have been driven to local maxima in the seven-dimensional parameter space by evolution, although dynamic range is near the maximum of what it could be. This implies that channel capacity is the subject of evolutionary selective pressure, adding to the literature on optimized properties of the bacterial chemotactic sensors.
@@ II.2 Monod-Wyman-Changeux Model of Bacterial Chemotactic Receptors
From Ref. [11], we have
@@
By substituting Eq. 2 into 1 and dividing the numerator and denominator by $e^{-\beta\varepsilon_{A}}\!\left(1+\frac{c}{K_{d}^{(A),1}}\right)^{N_{1}}\!\left(1+\frac{c}{K_{d}^{(A),2}}\right)^{N_{2}}$ yields
@@
Adaptation [19] is ignored not only for simplicity, but because in natural environments with large bacterial populations, chemoattractant and chemoattractant gradients constantly vanish [4]. Thus, negligible background concentration is assumed.
@@ II.3 Channel capacity
Channel capacity is the maximum mutual information that the receptor channel can transmit between the input ligand concentration and the output (activity). Mutual information $I(C;S)$ quantifies the average amount of information obtained about the input $c$ by observing the output $s\in\{I,A\}$. For each $c$, the conditional law is given by Eq. 3: $p(s{=}1\mid c)=p_{\mathrm{active}}(c)$ and $p(s{=}0\mid c)=1-p_{\mathrm{active}}(c)$. Let $p(c)$ be an arbitrary input distribution over concentrations and define the output marginal $p(s)=\int p(c)\,p(s\mid c)\,dc$. The mutual information in bits is given as
$I(C;S)=\sum_{s\in\{0,1\}}\int p(c)\,p(s\mid c)\,\log_{2}\!\frac{p(s\mid c)}{p(s)}\,dc.$ (4)
Since the channel has a binary output, $I(C;S)$ is between 0 and $1$ bits. $I(C;S)=0$ bits indicates that $c$ and $s$ are independent, while $I[C;S]=1$ bit indicates that the output is deterministic. Noise in the sensor or environment decreases mutual information. Channel capacity is the maximum mutual information obtained by choosing the most informative input distribution. Mathematically, it is the supremum over input distributions, given as
$C=\sup_{p(c)}I(C;S).$ (5)
Biologically, one must argue that the sensor changes its input concentration via movement so as to saturate its information processing capabilities.
@@ III Methods
To evaluate how well the receptor cluster can sense ligand concentration, its sensor properties can be described by three performance metrics: Channel capacity, effective hill coefficient, and dynamic range. We compute these metrics from the MWC activity curve $p_{\mathrm{active}}(c)$ and sweep seven model parameters: the allosteric constant $L_{0}$, the four state-dependent dissociation constants $K_{d}^{(I),1}$, $K_{d}^{(A),1}$, $K_{d}^{(I),2}$, $K_{d}^{(A),2}$, and the Tar/Tsr receptor copy numbers $N_{1}$ and $N_{2}$.
To visualize how these metrics depend on parameters, we plot two-dimensional slices through the seven-dimensional landscape as heatmaps. In each panel, the two plotted parameters span the axes, while the remaining five parameters are aggregated by taking the per-pixel median over all sampled values, yielding a representative summary of performance across the full sweep. Axes for $K_{d}^{(s),i}$ and $L_{0}$ are displayed on logarithmic scales, while axes for $N_{1}$ and $N_{2}$ are linear.
For visualization purposes, heatmap color scales are rescaled using percentile clipping, with limits set by the 2nd and 98th percentiles of the metric values within each panel. This procedure prevents extreme outliers from saturating the colormap and allows internal structure near biologically relevant regions to be resolved. All numerical analyses, correlations, and gradient calculations are performed on the unclipped data.
@@ III.1 Channel capacity
Channel capacity $C$ is numerically calculated via the Blahut–Arimoto algorithm [3, 1], which iteratively reweights the input distribution to monotonically increase $I(C;S)$ until convergence. In addition to the capacity value $C^{*}$, the algorithm also outputs the optimal Blahut–Arimoto input distribution $p^{*}(c)$ and the optimal output occupancy $p^{*}(s)$. $p^{*}(c)$ is the distribution of ligand concentrations that would realize that capacity, and $p^{*}(s)$ shows the fraction of time that the receptor ends up inactive versus active states at capacity.
We model receptor-cluster sensing as a discrete memoryless channel whose input is ligand concentration and whose output is the binary activity state of the cluster. The concentration axis is represented by a log-spaced grid $\{c_{j}\}_{j=1}^{M}$ of $M$ bins chosen to sweep over concentrations that showcase both the plateaus and the steep changes in activity curves of Fig. 3. Refining the log-spaced grid $\{c_{j}\}$ left qualitative conclusions unchanged. For each bin $c_{j}$, the output $s\in\{0,1\}$ (0 == inactive, 1 == active) is distributed according to the MWC response (Eq. 3):
$P(s=1\mid c_{j})$ $=p_{\mathrm{active}}(c_{j}),$ (6)
$P(s=0\mid c_{j})$ $=1-p_{\mathrm{active}}(c_{j}).$
Let $p(c_{j})$ denote the prior probability that the input lies in bin $j$. The induced output marginal is $p(s)=\sum_{j=1}^{M}p(c_{j})\,P(s\mid c_{j})$ for $s\in\{0,1\}$.
@@
To compute $C$ and a maximizing prior, we use the Blahut–Arimoto (BA) algorithm for fixed discrete channels. Starting from any initial prior $p^{(0)}(c_{j})$ with $\sum_{j}p^{(0)}(c_{j})=1$, define at iteration $t$ the output marginal
$p^{(t)}(s)=\sum_{j=1}^{M}p^{(t)}(c_{j})\,P(s\mid c_{j}),\qquad s\in\{0,1\}.$ (9)
@@
$p^{(t+1)}(c_{j})=\frac{\exp\!\big(\ell^{(t)}_{j}\big)}{\displaystyle\sum_{k=1}^{M}\exp\!\big(\ell^{(t)}_{k}\big)},\qquad j=1,\dots,M,$ (11)
which guarantees $\sum_{j}p^{(t+1)}(c_{j})=1$. These iterations monotonically increase $I(C;S)$ and converge to a capacity-achieving prior $p^{*}(c_{j})$. We report the capacity $C^{*}=I_{p^{*}}(C;S)$ (converted to bits via $\log_{2}$), together with the optimal input distribution $p^{*}(c_{j})$ and the induced optimal output occupancy
$p^{*}(s)=\sum_{j=1}^{M}p^{*}(c_{j})\,P(s\mid c_{j}),\qquad s\in\{0,1\}.$ (12)
Because the output is binary, the maximizing prior typically concentrates probability near concentrations whose outputs are most distinct.
@@ III.3 Effective Hill coefficient
Physically, the effective Hill coefficient $n_{\mathrm{eff}}$ quantifies how abruptly the output changes as $c$ crosses its midpoint, indicating that the binding sites act in concert (cooperativity). Negative cooperativity comes from $|n_{\mathrm{eff}}|<1$, while positive cooperativity comes from $|n_{\mathrm{eff}}|>1$. Greater deviations of $|n_{\mathrm{eff}}|$ from $1$ means a more switch-like response (small input changes cause large output changes near the concentration that leads to middling $p(c)$), but may narrow the dynamic range of the sensor.
@@
and define $c^{*}$ as the unique concentration that satisfy
@@
As shown in Supplementary Information,
$n_{\mathrm{eff}}$ $=$  $-\,4\,\frac{c^{*}\,p^{*}(1-p^{*})}{\,p_{\infty}-p_{0}\,}\,\Bigg[N_{2}\!\left(\frac{1}{K_{d}^{(I),1}+c^{*}}-\frac{1}{K_{d}^{(A),1}+c^{*}}\right)$ (21)
@@ IV Results
The static sensing performance of the mixed Tar/Tsr receptor cluster is evaluated by varying seven parameters: $L_{0}$, $K_{d}^{(I),1}$, $K_{d}^{(A),1}$, $K_{d}^{(I),2}$, $K_{d}^{(A),2}$, $N_{1}$, $N_{2}$. For each set of parameters, three performance metrics are computed to evaluate the sensor’s response: dynamic range $\mathrm{DR}$, effective hill coefficient $n_{\mathrm{eff}}$, and channel capacity $C$. Two-dimensional slices of this seven-dimensional parameter space are visualized in heatmaps, with color scales indicating the magnitude of each metric. To anchor the analysis in biological data, the receptor composition data of ten strains of E. coli are extracted from literature and overlayed on the heatmap as white symbols (Table 1). These strain markers provide reference points for where real bacteria sit in parameter space.
To quantify relationships among the three static metrics across the 7D sweep, we computed pairwise correlation coefficient $R$ (Table 2).
All three metrics are positively correlated as was true for populations of MWC receptors [9]. Channel capacity correlates strongly with dynamic range ($R=0.978$) and moderately with the effective Hill coefficient ($R=0.600$); effective Hill coefficient and dynamic range are also moderately correlated ($R=0.481$). These correlations support the idea that high information transfer typically requires both a sufficiently steep response and a broad input span. Despite the strong correlations, we claim that it is possible to distinguish which quantities are potential selection pressures in a manner described below.
@@
For clarity, this section displays a representative subset of heatmaps with strain reference points overlaid. See the supplementary information section for the full set of heatmaps across ten strains.

### v3

# Apparent Selection Pressure for Dynamic Range and Channel Capacity in Bacterial Chemotactic Sensors

@@ Abstract
Bacterial chemotactic sensing converts noisy chemical signals into running and tumbling. We analyze the static sensing limits of mixed Tar/Tsr chemoreceptor clusters in individual Escherichia coli cells using a heterogeneous Monod–Wyman–Changeux (MWC) model. By sweeping a seven-dimensional parameter space, we compute three sensing performance metrics—channel capacity, dynamic range, and effective Hill coefficient. Across E. coli–like parameter regimes, we consistently observe pronounced global maxima of channel capacity and global maxima of the related dynamic range, whereas the effective Hill coefficient does not exhibit comparable optimization. The capacity-achieving input distribution is bimodal, which implies that individual cells maximize information by sampling both low- and high-concentration regimes. Together, these results suggest that, at the individual-cell level, channel capacity and dynamic range may be selected for in E. coli receptor clusters.
@@ I Introduction
A model organism for studying this hypothesis is bacteria. Their chemotactic receptors, responsible for the movement of bacterial cells in response to environmental chemical stimuli, have been the subject of a number of beautiful models [2, 7, 8, 9]. In order to survive, E. coli cells must take in information about chemoattractant concentration [10] so as to move to maximize chemoattractant intake. Recent measurements in E. coli cells have demonstrated that they operate close to optimal information transfer for gradient-climbing [11]. A recent study showed that bacteria optimize predictive information transfer given resource constraints [12]. Furthermore, basic theoretical analyses say that populations of bacteria should maximize predictive information subject to constraints on resources [13].
However, no studies have addressed the channel capacities of individual bacteria. While some studies on E. coli chemotactic receptors focus on the population level and dynamical behavior of cells [3, 13], the present study provides an analysis of the chemotactic ability of a single E. coli cell. Specifically, the mixed Tar/Tsr chemoreceptor clusters within the E. coli cell are modeled with a heterogeneous Monod-Wyman-Changeux molecule [2] under the assumption of adaptation to negligible background chemoattractant concentration [14]. Three sensor performance metrics (channel capacity, dynamic range, and effective Hill coefficient) are obtained by sweeping a seven-dimensional parameter space. By locating literature-based strains in these seven-dimensional landscapes, we find that channel capacity appears to have been driven to a global maximum for all available strains. Dynamic range and effective Hill coefficient are positively correlated with channel capacity across the seven-dimensional parameter space, but upon closer examination, the effective Hill coefficient does not appear to have been driven to even a local maxima in the seven-dimensional parameter space by evolution. Dynamic range, too, appears to have been globally maximized, and is closely related to channel capacity. This implies that channel capacity and dynamic range are the subject of evolutionary selective pressure, adding to the literature on optimized properties of the bacterial chemotactic sensors.
@@ II.2 Monod-Wyman-Changeux Model of Bacterial Chemotactic Receptors
We define $p(c)$ as the probability that the cluster is in the active conformation at ligand concentration $c$. From Ref. [2], we have
@@
By substituting Eq. 2 into Eq. 1 and dividing the numerator and denominator by $e^{-\beta\varepsilon_{A}}\!\left(1+\frac{c}{K_{d}^{(A),1}}\right)^{N_{1}}\!\left(1+\frac{c}{K_{d}^{(A),2}}\right)^{N_{2}}$ yields
@@
Adaptation [7] is ignored not only for simplicity, but also because in natural environments with large bacterial populations, chemoattractant and chemoattractant gradients constantly vanish [14]. Thus, negligible background concentration is assumed.
@@ II.3 Channel capacity
Channel capacity is the maximum mutual information that the receptor channel can transmit between the input ligand concentration and the output (activity). Mutual information $I(C;S)$ quantifies the average amount of information obtained about the input $c$ by observing the output $s\in\{0,1\}$, where $s=0$ denotes inactive and $s=1$ denotes active. For each $c$, the conditional law is given by Eq. 3: $p(s{=}1\mid c)=p(c)$ and $p(s{=}0\mid c)=1-p(c)$. Let $p_{\mathrm{in}}(c)$ be an arbitrary input distribution over concentrations and define the output marginal $p(s)=\int p_{\mathrm{in}}(c)\,p(s\mid c)\,dc$. The mutual information (in bits) is given as
$I(C;S)=\sum_{s\in\{0,1\}}\int p_{\mathrm{in}}(c)\,p(s\mid c)\log_{2}\!\left(\frac{p(s\mid c)}{p(s)}\right)\,dc.$ (4)
Since the channel has a binary output, $I(C;S)$ is between 0 and $1$ bits. $I(C;S)=0$ bit indicates that $c$ and $s$ are independent, while $I(C;S)=1$ bit indicates that the output is deterministic. Noise in the sensor or environment decreases mutual information. Channel capacity is the maximum mutual information obtained by choosing the most informative input distribution. Mathematically, it is the supremum over input distributions, given as
$C=\sup_{p_{\mathrm{in}}(c)}I(C;S).$ (5)
Biologically, one must argue that the sensor changes its input concentration via movement in order to saturate its information processing capabilities.
@@ III Methods
To evaluate how well the receptor cluster can sense ligand concentration, its sensor properties are quantified by three performance metrics: channel capacity, dynamic range, and effective Hill coefficient. These metrics are computed from the MWC activity curve $p(c)$ while sweeping seven model parameters: the allosteric constant $L_{0}$, the four state-dependent dissociation constants $K_{d}^{(I),1}$, $K_{d}^{(A),1}$, $K_{d}^{(I),2}$, $K_{d}^{(A),2}$, and the Tar/Tsr receptor copy numbers $N_{1}$ and $N_{2}$.
For each strain, we perform a seven-dimensional parameter sweep by constructing a grid centered around a reference strain (Table LABEL:tab:1). Each parameter is varied independently around its reference value while all others are held fixed, with logarithmic spacing used for the allosteric and dissociation constants and linear spacing used for the receptor copy numbers. This procedure defines a discrete seven-dimensional hyper-rectangle surrounding the strain operating point, on which all three performance metrics are evaluated.
To analyze and present the results of the sweep, we extract lower-dimensional representations from the seven-dimensional data. Two-dimensional slices through parameter space are visualized as heatmaps, in which two parameters span the axes while the remaining five are fixed at the strain reference values (Table LABEL:tab:1). The ten strain operating points are overlaid on these heatmaps as reference markers to indicate where real bacteria sit in the parameter space. For visualization purposes, heatmap color scales are rescaled using percentile clipping (2nd–98th percentile); all numerical analyses are performed on the unclipped data. One-dimensional slices are additionally used to generate channel capacity, dynamic range, and effective Hill coefficient curves along individual parameter directions.
@@ III.1 Channel capacity
Channel capacity $C$ is numerically calculated via the Blahut–Arimoto algorithm [15, 16], which iteratively reweights the input distribution to monotonically increase $I(C;S)$ until convergence. In addition to the capacity value $C^{*}$, the algorithm also outputs the optimal input distribution $p_{\mathrm{in}}^{*}(c)$ and the optimal output occupancy $p^{*}(s)$. $p_{\mathrm{in}}^{*}(c)$ is the distribution of ligand concentrations that would realize that capacity, and $p^{*}(s)$ shows the fraction of time that the receptor ends up inactive versus active states at capacity.
We model receptor-cluster sensing as a discrete memoryless channel whose input is ligand concentration and whose output is the binary activity state of the cluster. The concentration axis is represented by a logarithmic grid $\{c_{j}\}_{j=1}^{M}$ of $M$ bins chosen to sweep over concentrations that showcase both the plateaus and the steep changes in activity curves of Fig. 3. Refining the logarithmic grid $\{c_{j}\}$ left qualitative conclusions unchanged. For each bin $c_{j}$, the output $s\in\{0,1\}$ is distributed according to the MWC response (Eq. 3):
$P(s=1\mid c_{j})$ $=p(c_{j}),$ (6)
$P(s=0\mid c_{j})$ $=1-p(c_{j}).$
Let $p_{\mathrm{in}}(c_{j})$ denote the prior probability that the input lies in bin $j$. The induced output marginal is $p(s)=\sum_{j=1}^{M}p_{\mathrm{in}}(c_{j})\,P(s\mid c_{j})$.
@@
To compute $C$ and a maximizing prior, we use the Blahut–Arimoto (BA) algorithm for fixed discrete channels. Starting from any initial prior $p_{\mathrm{in}}^{(0)}(c_{j})$ with $\sum_{j=1}^{M}p_{\mathrm{in}}^{(0)}(c_{j})=1$, define at iteration $t$ the output marginal
$p^{(t)}(s)=\sum_{j=1}^{M}p_{\mathrm{in}}^{(t)}(c_{j})\,P(s\mid c_{j}),\qquad s\in\{0,1\}.$ (9)
@@
$p_{\text{in}}^{(t+1)}(c_{j})=\frac{p_{\text{in}}^{(t)}(c_{j})\,\exp\!\big(\ell_{j}^{(t)}\big)}{\sum_{k}p_{\text{in}}^{(t)}(c_{k})\,\exp\!\big(\ell_{k}^{(t)}\big)}\,.$ (11)
which guarantees $\sum_{j}p_{\mathrm{in}}^{(t+1)}(c_{j})=1$. These iterations monotonically increase $I(C;S)$ and converge to a capacity-achieving prior $p_{\mathrm{in}}^{*}(c_{j})$. We report the capacity $C^{*}=I_{p_{\mathrm{in}}^{*}}(C;S)$ (converted to bits via $\log_{2}$), together with the optimal input distribution $p_{\mathrm{in}}^{*}(c_{j})$ and the induced optimal output occupancy
$p^{*}(s)=\sum_{j=1}^{M}p_{\mathrm{in}}^{*}(c_{j})\,P(s\mid c_{j}),\qquad s\in\{0,1\}.$ (12)
Because the output is binary, the maximizing prior typically concentrates probability near concentrations whose outputs are the most distinct.
@@ III.3 Effective Hill coefficient
Physically, the effective Hill coefficient $n_{\mathrm{eff}}$ quantifies how abruptly the output changes as $c$ crosses its midpoint, indicating that the binding sites act in concert (cooperativity). For attractants, the activity response $p(c)$ is decreasing; we report $\lvert n_{\mathrm{eff}}\rvert$ as the steepness magnitude. Negative cooperativity comes from $|n_{\mathrm{eff}}|<1$, while positive cooperativity comes from $|n_{\mathrm{eff}}|>1$. Greater deviations of $|n_{\mathrm{eff}}|$ from $1$ mean a more switch-like response (small input changes cause large output changes near the concentration that leads to middling $p(c)$), but may narrow the dynamic range of the sensor.
@@
and define $c^{*}$ as the unique concentration that satisfies
@@
As shown in the Supplementary Information Sec. S1,
$n_{\mathrm{eff}}$ $=$  $-\,4\,\frac{c^{*}\,p^{*}(1-p^{*})}{\,p_{\infty}-p_{0}\,}\,\Bigg[N_{1}\!\left(\frac{1}{K_{d}^{(I),1}+c^{*}}-\frac{1}{K_{d}^{(A),1}+c^{*}}\right)$ (21)
@@
III.4 Numerical gradient calculation
To interpret the seven-dimensional sweep in a way that is comparable across strains, we summarize the results using two complementary quantities. This assesses whether biological strains are positioned near globally optimal sensing performance and whether they lie in locally flat or steep regions of the performance landscape.
For each performance metric $M\in\{C,\mathrm{DR},n_{\mathrm{eff}}\}$, we quantify local flatness at the strain operating point by the Euclidean ($L_{2}$) norm of the gradient in a mixed coordinate system,
$\mathbf{u}$ $=$ $\bigl(\log_{10}L_{0},\log_{10}K_{d}^{(I),1},\log_{10}K_{d}^{(A),1},\log_{10}K_{d}^{(I),2}$ (22)
$,\log_{10}K_{d}^{(A),2},N_{1},N_{2}\bigr),$
where the five biochemical parameters are treated on a logarithmic scale and receptor copy numbers are treated linearly. At the strain parameter vector $\mathbf{u}$, partial derivatives are estimated numerically using fixed-step finite differences. For a logarithmic parameter $x\in\{L_{0},K_{d}^{(s),i}\}$, we approximate the derivative with respect to $\log_{10}x$ by
$\frac{\partial M}{\partial\log_{10}x}\approx\frac{M(x\,10^{\delta})-M(x\,10^{-\delta})}{2\delta},\qquad\delta=0.01,$ (23)
and for receptor copy numbers $N\in\{N_{1},N_{2}\}$ we use
$\frac{\partial M}{\partial N}\approx\frac{M(N+\Delta)-M(N-\Delta)}{2\Delta},\qquad\Delta=1.$ (24)
When a symmetric step leaves the swept bounds, we use the corresponding one-sided difference. The local gradient norm is then
$\|\nabla_{\mathbf{u}}M\|_{2}=\left(\sum_{i=1}^{7}\left(\frac{\partial M}{\partial u_{i}}\right)^{2}\right)^{1/2}.$ (25)
@@ IV Results
To quantify the relationships among the three static metrics in the 7D sweep, we computed pairwise correlation coefficients $R$ over all sampled grid points in the 7D sweep (Table II). All three metrics are positively correlated, as was true for populations of MWC receptors [3]. Channel capacity correlates strongly with dynamic range ($R=0.990$) and moderately with the effective Hill coefficient ($R=0.559$); effective Hill coefficient and dynamic range are also moderately correlated ($R=0.486$). These correlations support the idea that high information transfer typically requires both a sufficiently steep response and a broad input span. Despite the strong correlations, we claim that it is possible to distinguish which quantities are potential selection pressures in a manner described below.
@@
For clarity, this section displays a representative subset of heatmaps with strain reference points overlaid. See Supplementary Information Sec. S2 for the full set of heatmaps across ten strains.